# Recursive construction and enumeration of self-orthogonal and self-dual codes over finite commutative chain rings of even characteristic - I

### v1

@@ 2 Some preliminaries
      if $\ell$ is odd, where $\mathbf{v}_{a}^{(b)}$ denotes the $a$-th row of the matrix $\mathtt{V}^{(b)}]_{\gamma_{\ell}+1}$ for $1\leq a\leq\gamma_{\ell}+1-\mathtt{f}_{j+2}$ and $\ell\leq b\leq\mathtt{f}_{\ell+j-1+2\theta_{e}}.$ We next observe, by (3.3), that $\mathtt{E}^{(\ell+j)}_{a,a}=0$ for $1\leq a\leq\gamma_{\ell}+1-\mathtt{f}_{j+2},$ which implies that π_ℓ+j(Diag([W^(ℓ)]_γ_ℓ+1-f_j+2[W^(ℓ)]_γ_ℓ+1-f_j+2^t))=0 for $j\in\{\ell-1,\ell+1,\ldots,\kappa-2+\theta_{e}\}.$
    2. (ii)
      Next, let $\kappa\leq\ell\leq e-\kappa.$ Here, for $i\in\{2,4,\ldots,\kappa-1\}\cup\{\kappa\},$ as $\mathscr{D}_{\ell+i}$ is a self-orthogonal code over $\mathscr{R}_{\ell+i,m},$ by Lemma 2.1, we get $\mathtt{Diag}\big([\mathtt{W}^{(\ell+i)}]_{\gamma_{\ell}+1-\mathrm{c}_{i}}[\mathtt{W}^{(\ell+i)}]_{\gamma_{\ell}+1-\mathrm{c}_{i}}^{t}\big)\equiv\mathbf{0}\pmod{u^{\ell+i}}$, which implies that Diag([W^(ℓ)]_γ_ℓ+1-c_i[W^(ℓ)]_γ_ℓ+1-c_i^t)0(modu^ℓ+i).
    3. (iii)
      Finally, let $\ell\geq e-\kappa+1.$ Here, for $i\in\{2,4,\ldots,e-\ell\},$ since $\mathscr{D}_{\ell+i}$ is a self-orthogonal code over $\mathscr{R}_{\ell+i,m},$ by Lemma 2.1, we get $\mathtt{Diag}\big([\mathtt{W}^{(\ell+i)}]_{\gamma_{\ell}+1-\mathrm{f}_{i}}[\mathtt{W}^{(\ell+i)}]_{\gamma_{\ell}+1-\mathrm{f}_{i}}^{t}\big)\equiv\mathbf{0}\pmod{u^{\ell+i}}$, which implies that Diag([W^(ℓ)]_γ_ℓ+1-f_i[W^(ℓ)]_γ_ℓ+1-f_i^t)0(modu^ℓ+i).
@@
From the above lemma, we see, for $2\leq\ell\leq e-2$ and $\ell\equiv e\pmod{2}$, that a necessary condition for lifting a linear code $\mathscr{D}_{\ell}$ over $\mathscr{R}_{\ell,m}$ to a self-orthogonal (resp. self-dual) code $\mathscr{D}_{e}$ over $\mathscr{R}_{e,m}$ is that $\mathscr{D}_{\ell}$ must be a self-orthogonal (resp. self-dual) code over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P}).$ We next prove the following lemma.
@@
Further, let $\mathtt{Sym}_{j}(\mathcal{T}_{m})$ and $\mathtt{Alt}_{j}(\mathcal{T}_{m})$ denote the sets consisting of all $j\times j$ symmetric and alternating matrices over $\mathcal{T}_{m},$ respectively. We further recall, from Section 2, that $2\equiv u^{\kappa}(\eta_{0}+u\eta_{1}+u^{2}\eta_{2}+\cdots+u^{e-1-\kappa}\eta_{e-1-\kappa})\pmod{u^{e}}$, where $\eta_{0}\in\mathcal{T}_{m}\setminus\{0\}$ and $\eta_{1},\eta_{2},\ldots,\eta_{e-1-\kappa}\in\mathcal{T}_{m}.$ Now, the following proposition considers the case where $e\geq 4$ is an even integer and $\kappa$ is an odd integer satisfying $3\leq\kappa\leq e-1.$ It provides a method for lifting a $\Lambda_{s-\kappa_{1}}$-doubly even code of length $n$ and dimension $\Lambda_{s}$ over $\mathcal{T}_{m}$ to a self-orthogonal code of type $\{\Lambda_{s},\lambda_{s+1}\}$ and the same length $n$ over $\mathscr{R}_{2,m}$ satisfying property $(\mathfrak{P})$ with also enumerates the number of possible ways to perform this lifting.
@@
    $(2^{m})^{\sum\limits_{i=3}^{s+1}\lambda_{i}\Lambda_{i-2}+\Lambda_{s}(n-\Lambda_{s+1})-\Lambda_{s-1}-\frac{\Lambda_{s}(\Lambda_{s}-1)}{2}}{\lambda_{s+1}+n-\Lambda_{s+1}-\Lambda_{s}\brack\lambda_{s+1}}_{2^{m}}$
@@
The following proposition considers the case when both $e$ and $\kappa$ are odd integers satisfying $3\leq\kappa\leq e-1.$ It provides a method for lifting a $\Lambda_{s-\kappa_{1}}$-doubly even code of length $n$ and dimension $\Lambda_{s+1}$ over $\mathcal{T}_{m}$ to a self-orthogonal code of type $\{\Lambda_{s},\lambda_{s+1},\lambda_{s+2}\}$ and the same length $n$ over $\mathscr{R}_{3,m}$ satisfying property $(\mathfrak{P})$ with also enumerates the number of possible ways to perform this lifting.
@@
$\mathtt{G}_{3}=\begin{bmatrix}\mathtt{W}_{1}^{(3)}\\ \mathtt{W}_{2}^{(3)}\\ \vdots\\ \mathtt{W}_{s}^{(3)}\\ u\mathtt{W}_{s+1}^{(3)}\\ u^{2}\mathtt{W}_{s+2}^{(3)}\end{bmatrix}=\begin{bmatrix}\mathtt{W}_{1}^{(0)}+u\mathtt{V}_{1}^{(1)}+u^{2}\mathtt{V}_{1}^{(2)}\\ \mathtt{W}_{2}^{(0)}+u\mathtt{V}_{2}^{(1)}+u^{2}\mathtt{V}_{2}^{(2)}\\ \vdots\\ \mathtt{W}_{s}^{(0)}+u\mathtt{V}_{s}^{(1)}+u^{2}\mathtt{V}_{s}^{(2)}\\ u\mathtt{W}_{s+1}^{(3)}\vskip 1.42262pt\\ u^{2}\mathtt{W}_{s+2}^{(3)}\end{bmatrix}$,
@@
First of all, we will establish the existence of the matrices $\mathtt{V}^{(1)}]_{s}$ and $\mathtt{V}^{(2)}]_{s}$ satisfying the system of matrix equations (3.4)-(3.6) and count their choices. To do this, let us suppose that $\mathtt{W}^{(0)}]_{s}=(\textbf{b}_{i}), [$$\mathtt{V}^{(1)}]_{s}=(\textbf{x}_{j})$ and $\mathtt{V}^{(2)}]_{s}=(\textbf{y}_{g}),$ where $\textbf{b}_{i}$’s, $\textbf{x}_{j}$’s, and $\textbf{y}_{g}$’s denote the rows of the matrices $\mathtt{W}^{(0)}]_{s}, [$$\mathtt{V}^{(1)}]_{s},$ and $\mathtt{V}^{(2)}]_{s},$ respectively. We further observe, for $1\leq i\leq\Lambda_{s},$ that the $(i,i)$-th entry, say $J_{i}$, of the matrix $\mathtt{Diag}\left([\mathtt{W}^{(3)}]_{s}[\mathtt{W}^{(3)}]_{s}^{t}\right)$ satisfies
@@
and hence there exists a matrix $\textbf{[}\mathtt{V}^{(1)}\textbf{]}_{s}\in\mathcal{M}_{\Lambda_{s}\times n}(\mathcal{T}_{m})$ satisfying (3.13) and (3.14). Moreover, such a matrix $\textbf{[}\mathtt{V}^{(1)}\textbf{]}_{s}\in\mathcal{M}_{\Lambda_{s}\times n}(\mathcal{T}_{m})$ has precisely
$(2^{m})^{\sum\limits_{i=3}^{s+2}\lambda_{i}\Lambda_{i-2}+\Lambda_{s}(n-\Lambda_{s+2})-\Lambda_{s}-\frac{\Lambda_{s}(\Lambda_{s}-1)}{2}}$
distinct choices. Next, for a given choice of the matrix $\mathtt{V}^{(1)}]_{s}$ satisfying (3.13) and (3.14), we can write
@@
$\left.+\eta_{0}\left([\mathtt{W}^{(0)}]_{\gamma_{\ell}-\kappa_{1}}[\mathtt{V}^{(\ell-1)}]_{\gamma_{\ell}-\kappa_{1}}^{t}+[\mathtt{V}^{(1)}]_{\gamma_{\ell}-\kappa_{1}}[\mathtt{V}^{(\ell-2)}]_{\gamma_{\ell}-\kappa_{1}}^{t}\right)\right)$ $\equiv$ $\mathbf{0}\pmod{u}\text{ if }\theta_{e}=1,$~~~~~~~ (3.20)
$\mathtt{Diag}\left(\mathtt{E}^{(\ell+\kappa-2)}+\eta_{0}[\mathtt{W}^{(0)}]_{\gamma_{\ell}+1-\kappa_{1}}[\mathtt{V}^{(\ell-2)}]_{\gamma_{\ell}+1-\kappa_{1}}^{t}\right)$ $\equiv$ $\mathbf{0}\pmod{u}\text{ if }\theta_{e}=0,$~~~~~~~ (3.21)
@@
First of all, we will establish the existence of the matrices $\mathtt{V}^{(\ell-2)}]_{\gamma_{\ell}+1}$ and $\mathtt{V}^{(\ell-1)}]_{\gamma_{\ell}+1}$ satisfying the system of matrix equations (3.17)-(3.21). To do this, we will distinguish the following two cases: (I) $\ell$ is even, and (II) $\ell$ is odd.
@@
    Towards this, let $\mathtt{W}^{(0)}]_{\gamma_{\ell}+1}=(\textbf{b}_{i})$ and $\mathtt{V}^{(\ell-2)}]_{\gamma_{\ell}+1}=(\textbf{v}_{j}),$ where $\textbf{b}_{i}$ and $\textbf{v}_{j}$ denote the $i$-th and $j$-th rows of the matrices $\mathtt{W}^{(0)}]_{\gamma_{\ell}+1}$ and $\mathtt{V}^{(\ell-2)}]_{\gamma_{\ell}+1},$ respectively, for each $i$ and $j.$ Further, let $\mathtt{F}^{(\ell-2)}_{i,j},\mathtt{E}^{(2\ell-4)}_{i,j},\mathtt{E}^{(\ell+\kappa-2)}_{i,j}\in\mathcal{T}_{m}$ denote the $(i,j)$-th entries of the matrices $\mathtt{F}^{(\ell-2)},\mathtt{E}^{(2\ell-4)},$ $\mathtt{E}^{(\ell+\kappa-2)},$ respectively. Thus, the system of matrix equations (3.24)-(3.26) is equivalent to the following system of equations in unknowns $\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{\gamma_{\ell}+1}}$ over $\mathcal{T}_{m}$:
@@
    For each integer $j$ satisfying $1\leq j\leq\Lambda_{\gamma_{\ell}+1},$ one can easily see that there exists a unique integer $r_{j}$ satisfying $1\leq r_{j}\leq\gamma_{\ell}+1$ and $\Lambda_{r_{j}-1}+1\leq j\leq\Lambda_{r_{j}}.$ The corresponding unknown vector $\textbf{v}_{j}$ can be written as $\textbf{v}_{j}=(\textbf{0}~\textbf{v}_{j}^{n-\Lambda_{r_{j}+\ell-2}}),$ where 0 denotes the zero vector of length $\Lambda_{r_{j}+\ell-2}$ and $\textbf{v}_{j}^{n-\Lambda_{r_{j}+\ell-2}}$ denotes the vector of length $n-\Lambda_{r_{j}+\ell-2}$ obtained by omitting the first $\Lambda_{r_{j}+\ell-2}$ coordinates of $\textbf{v}_{j}$. This implies, for $\Lambda_{r_{j}-1}+1\leq j\leq\Lambda_{r_{j}}$, that the first $\Lambda_{r_{j}+\ell-2}$ coordinates of $\textbf{v}_{j}$ are zero, leaving $n-\Lambda_{r_{j}+\ell-2}$ variables in $\textbf{v}_{j}.$ For $1\leq j\leq\Lambda_{\gamma_{\ell}+1},$ let $\widehat{\textbf{v}}_{j}=\textbf{v}_{j}^{n-\Lambda_{r_{j}+\ell-2}}$ and $\widehat{\textbf{b}}_{j}=\textbf{b}_{j}^{n-\Lambda_{r_{j}+\ell-2}}$ be the vectors of length $n-\Lambda_{r_{j}+\ell-2}$ obtained from $\textbf{v}_{j}$ and $\textbf{b}_{j}$ by omitting their first $\Lambda_{r_{j}+\ell-2}$ coordinates, respectively. Consequently, the system of equations (3.27)-(3.29) is equivalent to the following system of equations in unknowns $\widehat{\textbf{v}}_{1},\widehat{\textbf{v}}_{2},\ldots,\widehat{\textbf{v}}_{\Lambda_{\gamma_{\ell}+1}}$ over $\mathcal{T}_{m}$:
@@
    where $\widehat{\textbf{1}}$ denotes the all-one vector having the same length as that of the vector $\widehat{\textbf{v}}_{i}$ for each $i.$ We further express the above system of equations (3.30)-(3.32) in the matrix form as
@@
      $(2^{m})^{\sum\limits_{i=\ell}^{\gamma_{\ell}+\ell}\lambda_{i}\Lambda_{i-\ell+1}+\Lambda_{\gamma_{\ell}+1}(n-\Lambda_{\gamma_{\ell}+\ell})-\Lambda_{s-2\mathrm{f}_{\ell}+2}-\Lambda_{\gamma_{\ell}+1-\kappa_{1}}-\frac{\Lambda_{\gamma_{\ell}+1}(\Lambda_{\gamma_{\ell}+1}-1)}{2}}.$
@@
      $\sum\limits_{h=1}^{\Lambda_{\gamma_{\ell}+1-\kappa_{1}}}\alpha_{h}(\mathtt{E}^{(2\ell-4)}_{h,h})^{2^{m-1}}+\eta_{0}^{-1}\sum\limits_{g=1}^{\Lambda_{\gamma_{\ell}+1-\kappa_{1}}}\alpha_{g}^{2}\mathtt{E}^{(\ell+\kappa-2)}_{g,g}+\hskip-11.38109pt\sum\limits_{~~~~~1\leq i<j\leq\Lambda_{\gamma_{\ell}+1-\kappa_{1}}}\hskip-11.38109pt\alpha_{i}\alpha_{j}\mathtt{F}_{i,j}^{(\ell-2)}~\equiv~0\pmod{u}$. (3.34)
@@
      $\textbf{1}\equiv\sum\limits_{j=1}^{\Lambda_{\gamma_{\ell}+1-\kappa_{1}}}\hskip-5.69054pt\alpha_{j}\left(\textbf{b}_{j}+u\textbf{v}^{(1)}_{j}+\cdots+u^{\ell-3}\textbf{v}^{(\ell-3)}_{j}\right)+u^{\ell-2}\textbf{d}_{\ell-2}+\cdots+u^{\ell+\kappa-2}\textbf{d}_{\ell+\kappa-2}\pmod{u^{\ell+\kappa-1}}$
@@
      $\sum\limits_{h=1}^{\Lambda_{\gamma_{\ell}+1-\kappa_{1}}}\alpha_{h}^{2}\mathtt{E}_{h,h}^{(2\ell-4)}+\textbf{d}_{\ell-2}\cdot\textbf{d}_{\ell-2}\equiv 0\pmod{u}$ (3.35)
@@
      $\eta_{0}^{-1}\sum\limits_{g=1}^{\Lambda_{\gamma_{\ell}+1-\kappa_{1}}}\alpha_{g}^{2}\mathtt{E}^{(\ell+\kappa-2)}_{g,g}+\hskip-14.22636pt\sum\limits_{~~~~~1\leq i<j\leq\Lambda_{\gamma_{\ell}+1-\kappa_{1}}}\hskip-14.22636pt\alpha_{i}\alpha_{j}\mathtt{F}_{i,j}^{(\ell-2)}+\textbf{1}\cdot\textbf{d}_{\ell-2}\equiv 0\pmod{u}$. (3.36)

### v2

@@ 2.6 Notations and preliminary setup
if $\ell$ is odd, where $\mathbf{v}_{a}^{(b)}$ denotes the $a$-th row of the matrix $\mathtt{V}^{(b)}]_{\gamma_{\ell}+1}$ for $1\leq a\leq\gamma_{\ell}+1-\mathtt{f}_{j+2}$ and $\ell\leq b\leq\mathtt{f}_{\ell+j-1+2\theta_{e}}.$ We next observe, by (3.3), that $\mathtt{E}^{(\ell+j)}_{a,a}=0$ for $1\leq a\leq\gamma_{\ell}+1-\mathtt{f}_{j+2},$ which implies that
$\pi_{\ell+j}\big(\mathtt{Diag}\big([\mathtt{W}^{(\ell)}]_{\gamma_{\ell}+1-\mathrm{f}_{j+2}}[\mathtt{W}^{(\ell)}]_{\gamma_{\ell}+1-\mathrm{f}_{j+2}}^{t}\big)\big)=\mathbf{0}$
for $j\in\{\ell-1,\ell+1,\ldots,\kappa-2+\theta_{e}\}.$
Next, let $\kappa\leq\ell\leq e-\kappa.$ Here, for $i\in\{2,4,\ldots,\kappa-1\}\cup\{\kappa\},$ as $\mathscr{D}_{\ell+i}$ is a self-orthogonal code over $\mathscr{R}_{\ell+i,m},$ by Lemma 2.1, we get $\mathtt{Diag}\big([\mathtt{W}^{(\ell+i)}]_{\gamma_{\ell}+1-\mathrm{c}_{i}}[\mathtt{W}^{(\ell+i)}]_{\gamma_{\ell}+1-\mathrm{c}_{i}}^{t}\big)\equiv\mathbf{0}\pmod{u^{\ell+i}}$, which implies that
$\mathtt{Diag}\big([\mathtt{W}^{(\ell)}]_{\gamma_{\ell}+1-\mathrm{c}_{i}}[\mathtt{W}^{(\ell)}]_{\gamma_{\ell}+1-\mathrm{c}_{i}}^{t}\big)\equiv\mathbf{0}\pmod{u^{\ell+i}}$.
Finally, let $\ell\geq e-\kappa+1.$ Here, for $i\in\{2,4,\ldots,e-\ell\},$ since $\mathscr{D}_{\ell+i}$ is a self-orthogonal code over $\mathscr{R}_{\ell+i,m},$ by Lemma 2.1, we get $\mathtt{Diag}\big([\mathtt{W}^{(\ell+i)}]_{\gamma_{\ell}+1-\mathrm{f}_{i}}[\mathtt{W}^{(\ell+i)}]_{\gamma_{\ell}+1-\mathrm{f}_{i}}^{t}\big)\equiv\mathbf{0}\pmod{u^{\ell+i}}$, which implies that
$\mathtt{Diag}\big([\mathtt{W}^{(\ell)}]_{\gamma_{\ell}+1-\mathrm{f}_{i}}[\mathtt{W}^{(\ell)}]_{\gamma_{\ell}+1-\mathrm{f}_{i}}^{t}\big)\equiv\mathbf{0}\pmod{u^{\ell+i}}$.
@@
From the above lemma, we see, for $2\leq\ell\leq e-2$ and $\ell\equiv e\pmod{2}$, that a necessary condition for lifting a linear code $\mathscr{D}_{\ell}$ over $\mathscr{R}_{\ell,m}$ to a self-orthogonal (resp. self-dual) code $\mathscr{D}_{e}$ over $\mathscr{R}_{e,m}$ is that $\mathscr{D}_{\ell}$ itself must be a self-orthogonal (resp. self-dual) code over $\mathscr{R}_{\ell,m}$ and must additionally satisfy property $(\mathfrak{P}).$ We next prove the following lemma.
@@
Further, let $\mathtt{Sym}_{j}(\mathcal{T}_{m})$ and $\mathtt{Alt}_{j}(\mathcal{T}_{m})$ denote the sets consisting of all $j\times j$ symmetric and alternating matrices over $\mathcal{T}_{m},$ respectively. We further recall, from Section 2, that $2\equiv u^{\kappa}(\eta_{0}+u\eta_{1}+u^{2}\eta_{2}+\cdots+u^{e-1-\kappa}\eta_{e-1-\kappa})\pmod{u^{e}}$, where $\eta_{0}\in\mathcal{T}_{m}\setminus\{0\}$ and $\eta_{1},\eta_{2},\ldots,\eta_{e-1-\kappa}\in\mathcal{T}_{m}.$ Now, the following proposition considers the case where $e\geq 4$ is an even integer and $\kappa$ is an odd integer satisfying $3\leq\kappa\leq e-1.$ It provides a method for lifting a $\Lambda_{s-\kappa_{1}}$-doubly even code of length $n$ and dimension $\Lambda_{s}$ over $\mathcal{T}_{m}$ to a self-orthogonal code of type $\{\Lambda_{s},\lambda_{s+1}\}$ and the same length $n$ over $\mathscr{R}_{2,m}$ satisfying property $(\mathfrak{P}),$ and it also enumerates the number of distinct ways to perform this lifting.
@@
    $(2^{m})^{\sum\limits_{i=3}^{s+1}\lambda_{i}\Lambda_{i-2}+\Lambda_{s}(n-\Lambda_{s+1})-\Lambda_{s-1}-\frac{\Lambda_{s}(\Lambda_{s}-1)}{2}}{\lambda_{s+1}+n-\Lambda_{s+1}-\Lambda_{s}\brack\lambda_{s+1}}_{2^{m}}\vskip-2.84526pt$
@@
The following proposition considers the case when both $e$ and $\kappa$ are odd integers satisfying $3\leq\kappa\leq e-1.$ It provides a method for lifting a $\Lambda_{s-\kappa_{1}}$-doubly even code of length $n$ and dimension $\Lambda_{s+1}$ over $\mathcal{T}_{m}$ to a self-orthogonal code of type $\{\Lambda_{s},\lambda_{s+1},\lambda_{s+2}\}$ and the same length $n$ over $\mathscr{R}_{3,m}$ satisfying property $(\mathfrak{P}),$ and it also enumerates the number of distinct ways in which this lifting can be performed.
@@
$\vskip-2.84526pt\mathtt{G}_{3}=\begin{bmatrix}\mathtt{W}_{1}^{(3)}\\ \mathtt{W}_{2}^{(3)}\\ \vdots\\ \mathtt{W}_{s}^{(3)}\\ u\mathtt{W}_{s+1}^{(3)}\\ u^{2}\mathtt{W}_{s+2}^{(3)}\end{bmatrix}=\begin{bmatrix}\mathtt{W}_{1}^{(0)}+u\mathtt{V}_{1}^{(1)}+u^{2}\mathtt{V}_{1}^{(2)}\\ \mathtt{W}_{2}^{(0)}+u\mathtt{V}_{2}^{(1)}+u^{2}\mathtt{V}_{2}^{(2)}\\ \vdots\\ \mathtt{W}_{s}^{(0)}+u\mathtt{V}_{s}^{(1)}+u^{2}\mathtt{V}_{s}^{(2)}\\ u\mathtt{W}_{s+1}^{(3)}\vskip 1.42262pt\\ u^{2}\mathtt{W}_{s+2}^{(3)}\end{bmatrix}$,
@@
First of all, we will establish the existence of the matrices $\mathtt{V}^{(1)}]_{s}$ and $\mathtt{V}^{(2)}]_{s}$ satisfying the system of matrix equations (3.4)–(3.6) and count their choices. To do this, let us suppose that $\mathtt{W}^{(0)}]_{s}=(\textbf{b}_{i}), [$$\mathtt{V}^{(1)}]_{s}=(\textbf{x}_{j})$ and $\mathtt{V}^{(2)}]_{s}=(\textbf{y}_{g}),$ where $\textbf{b}_{i}$’s, $\textbf{x}_{j}$’s, and $\textbf{y}_{g}$’s denote the rows of the matrices $\mathtt{W}^{(0)}]_{s}, [$$\mathtt{V}^{(1)}]_{s},$ and $\mathtt{V}^{(2)}]_{s},$ respectively. We further observe, for $1\leq i\leq\Lambda_{s},$ that the $(i,i)$-th entry, say $J_{i}$, of the matrix $\mathtt{Diag}\left([\mathtt{W}^{(3)}]_{s}[\mathtt{W}^{(3)}]_{s}^{t}\right)$ satisfies
@@
and hence there exists a matrix $\textbf{[}\mathtt{V}^{(1)}\textbf{]}_{s}\in\mathcal{M}_{\Lambda_{s}\times n}(\mathcal{T}_{m})$ satisfying (3.13) and (3.14). Moreover, such a matrix $\textbf{[}\mathtt{V}^{(1)}\textbf{]}_{s}\in\mathcal{M}_{\Lambda_{s}\times n}(\mathcal{T}_{m})$ has precisely $(2^{m})^{\sum\limits_{i=3}^{s+2}\lambda_{i}\Lambda_{i-2}+\Lambda_{s}(n-\Lambda_{s+2})-\Lambda_{s}-\frac{\Lambda_{s}(\Lambda_{s}-1)}{2}}$ distinct choices. Next, for a given choice of the matrix $\mathtt{V}^{(1)}]_{s}$ satisfying (3.13) and (3.14), we can write
@@
$\left.+\eta_{0}\left([\mathtt{W}^{(0)}]_{\gamma_{\ell}-\kappa_{1}}[\mathtt{V}^{(\ell-1)}]_{\gamma_{\ell}-\kappa_{1}}^{t}+[\mathtt{V}^{(1)}]_{\gamma_{\ell}-\kappa_{1}}[\mathtt{V}^{(\ell-2)}]_{\gamma_{\ell}-\kappa_{1}}^{t}\right)\right)$ $\equiv$ $\mathbf{0}\pmod{u}\text{ if }e\text{ is odd},$~~~~~~~ (3.20)
$\mathtt{Diag}\left(\mathtt{E}^{(\ell+\kappa-2)}+\eta_{0}[\mathtt{W}^{(0)}]_{\gamma_{\ell}+1-\kappa_{1}}[\mathtt{V}^{(\ell-2)}]_{\gamma_{\ell}+1-\kappa_{1}}^{t}\right)$ $\equiv$ $\mathbf{0}\pmod{u}\text{ if }e\text{ is even},$~~~~~~~ (3.21)
@@
First of all, we will establish the existence of the matrices $\mathtt{V}^{(\ell-2)}]_{\gamma_{\ell}+1}$ and $\mathtt{V}^{(\ell-1)}]_{\gamma_{\ell}+1}$ satisfying the system of matrix equations (3.17)–(3.21). To do this, we will distinguish the following two cases: (I) $\ell$ is even, and (II) $\ell$ is odd.
@@
    Towards this, let $\mathtt{W}^{(0)}]_{\gamma_{\ell}+1}=(\textbf{b}_{i})$ and $\mathtt{V}^{(\ell-2)}]_{\gamma_{\ell}+1}=(\textbf{v}_{j}),$ where $\textbf{b}_{i}$ and $\textbf{v}_{j}$ denote the $i$-th and $j$-th rows of the matrices $\mathtt{W}^{(0)}]_{\gamma_{\ell}+1}$ and $\mathtt{V}^{(\ell-2)}]_{\gamma_{\ell}+1},$ respectively, for each $i$ and $j.$ Further, let $\mathtt{F}^{(\ell-2)}_{i,j},\mathtt{E}^{(2\ell-4)}_{i,j},\mathtt{E}^{(\ell+\kappa-2)}_{i,j}\in\mathcal{T}_{m}$ denote the $(i,j)$-th entries of the matrices $\mathtt{F}^{(\ell-2)},\mathtt{E}^{(2\ell-4)},$ $\mathtt{E}^{(\ell+\kappa-2)},$ respectively. Thus, the system of matrix equations (3.24)–(3.26) is equivalent to the following system of equations in unknowns $\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{\gamma_{\ell}+1}}$ over $\mathcal{T}_{m}$:
@@
    For each integer $j$ satisfying $1\leq j\leq\Lambda_{\gamma_{\ell}+1},$ one can easily see that there exists a unique integer $r_{j}$ satisfying $1\leq r_{j}\leq\gamma_{\ell}+1$ and $\Lambda_{r_{j}-1}+1\leq j\leq\Lambda_{r_{j}}.$ The corresponding unknown vector $\textbf{v}_{j}$ can be written as $\textbf{v}_{j}=(\textbf{0}~\textbf{v}_{j}^{n-\Lambda_{r_{j}+\ell-2}}),$ where 0 denotes the zero vector of length $\Lambda_{r_{j}+\ell-2}$ and $\textbf{v}_{j}^{n-\Lambda_{r_{j}+\ell-2}}$ denotes the vector of length $n-\Lambda_{r_{j}+\ell-2}$ obtained by omitting the first $\Lambda_{r_{j}+\ell-2}$ coordinates of $\textbf{v}_{j}$. This implies, for $\Lambda_{r_{j}-1}+1\leq j\leq\Lambda_{r_{j}}$, that the first $\Lambda_{r_{j}+\ell-2}$ coordinates of $\textbf{v}_{j}$ are zero, leaving $n-\Lambda_{r_{j}+\ell-2}$ variables in $\textbf{v}_{j}.$ For $1\leq j\leq\Lambda_{\gamma_{\ell}+1},$ let $\widehat{\textbf{v}}_{j}=\textbf{v}_{j}^{n-\Lambda_{r_{j}+\ell-2}}$ and $\widehat{\textbf{b}}_{j}=\textbf{b}_{j}^{n-\Lambda_{r_{j}+\ell-2}}$ be the vectors of length $n-\Lambda_{r_{j}+\ell-2}$ obtained from $\textbf{v}_{j}$ and $\textbf{b}_{j}$ by omitting their first $\Lambda_{r_{j}+\ell-2}$ coordinates, respectively. Consequently, the system of equations (3.27)–(3.29) is equivalent to the following system of equations in unknowns $\widehat{\textbf{v}}_{1},\widehat{\textbf{v}}_{2},\ldots,\widehat{\textbf{v}}_{\Lambda_{\gamma_{\ell}+1}}$ over $\mathcal{T}_{m}$:
@@
    where $\widehat{\textbf{1}}$ denotes the all-one vector having the same length as that of the vector $\widehat{\textbf{v}}_{i}$ for each $i.$ We further express the above system of equations (3.30)–(3.32) in the matrix form as
@@
      $(2^{m})^{\sum\limits_{i=\ell}^{\gamma_{\ell}+\ell}\lambda_{i}\Lambda_{i-\ell+1}+\Lambda_{\gamma_{\ell}+1}(n-\Lambda_{\gamma_{\ell}+\ell})-\Lambda_{s-2\mathrm{f}_{\ell}+2}-\Lambda_{\gamma_{\ell}+1-\kappa_{1}}-\frac{\Lambda_{\gamma_{\ell}+1}(\Lambda_{\gamma_{\ell}+1}-1)}{2}}.\vskip-5.69054pt$
@@
      $\vskip-2.84526pt\displaystyle\sum\limits_{h=1}^{\Lambda_{\gamma_{\ell}+1-\kappa_{1}}}\alpha_{h}(\mathtt{E}^{(2\ell-4)}_{h,h})^{2^{m-1}}+\eta_{0}^{-1}\sum\limits_{g=1}^{\Lambda_{\gamma_{\ell}+1-\kappa_{1}}}\alpha_{g}^{2}\mathtt{E}^{(\ell+\kappa-2)}_{g,g}+\hskip-11.38109pt\sum\limits_{~~~~~1\leq i<j\leq\Lambda_{\gamma_{\ell}+1-\kappa_{1}}}\hskip-11.38109pt\alpha_{i}\alpha_{j}\mathtt{F}_{i,j}^{(\ell-2)}~\equiv~0\pmod{u}$. (3.34)
@@
      $\vskip-2.84526pt\textbf{1}\equiv\sum\limits_{j=1}^{\Lambda_{\gamma_{\ell}+1-\kappa_{1}}}\hskip-5.69054pt\alpha_{j}\left(\textbf{b}_{j}+u\textbf{v}^{(1)}_{j}+\cdots+u^{\ell-3}\textbf{v}^{(\ell-3)}_{j}\right)+u^{\ell-2}\textbf{d}_{\ell-2}+\cdots+u^{\ell+\kappa-2}\textbf{d}_{\ell+\kappa-2}\pmod{u^{\ell+\kappa-1}}$
@@
      $\vskip-2.84526pt\sum\limits_{h=1}^{\Lambda_{\gamma_{\ell}+1-\kappa_{1}}}\alpha_{h}^{2}\mathtt{E}_{h,h}^{(2\ell-4)}+\textbf{d}_{\ell-2}\cdot\textbf{d}_{\ell-2}\equiv 0\pmod{u}$ (3.35)
@@
      $\vskip-2.84526pt\eta_{0}^{-1}\sum\limits_{g=1}^{\Lambda_{\gamma_{\ell}+1-\kappa_{1}}}\alpha_{g}^{2}\mathtt{E}^{(\ell+\kappa-2)}_{g,g}+\hskip-14.22636pt\sum\limits_{~~~~~1\leq i<j\leq\Lambda_{\gamma_{\ell}+1-\kappa_{1}}}\hskip-14.22636pt\alpha_{i}\alpha_{j}\mathtt{F}_{i,j}^{(\ell-2)}+\textbf{1}\cdot\textbf{d}_{\ell-2}\equiv 0\pmod{u}$. (3.36)